Manganese Company makes frames. A customer wants to place a special order for 750 frames in green with the company logo painted on the frame, to be priced at $60 each. Normally, Manganese would charge $100 per frame for this type of order. Manganese figures that wood and glass will cost $20 per frame, variable overhead (machining, electricity) is $5 per frame, direct labor is $10 per frame, nd one setup will be required at $1,500 per setup. The set-up charge costs are 100% labor. Currently, the workers needed to set up for and make the frames are working at Manganese. Their wages will be paid whether or not the special order is accepted. Manganeses policy is to avoid layoffs to the extent possible. If Manganese accepts the special order, by how much will operating income increase or decrease? Oa. There will be no effect on operating income. Ob. $40,000 decrease Oc. $26,250 increase Od. $30,000 increase Ce. $16,660 increase Next Previous
Colortrigon Company makes a variety of paper products. One product is 30 lb copler paper, packaged 3,000 sheets to a box. One box normally sells for $20. A large bank offered to purchase 6,000 boxes at $15 per box. Costs per box are as follows: Direct materials Direct labor $6 Fxed overhead No variable marketing costs would be incurred on the order. The company is operating significantly below the maximum productive capacity. No fixed costs are avoidable. Should Colortrigon accept the order? Oa. Yes, income will increase by $19,000 Ob. Yes, income will increase by $30,000. Oc. No, income will decrease by $86,000. Od. No, income will decrease by $43,000. Oe. It doesnt matter; there will be no impact on income. Next
Answer #1

1. Answer C. $26,250 increase

Total order 750 * 60 = 45000

Total cost = 20 + 5 = 25 * 750 = 18750

Operating income = Total revenue - direct cost - indirect cost

= 45000 - 18750 = 26250

Note: the wages of the labor will be paid whether or not the special order is accepted.

2. Answer B. Yes, income increase by $30,000

Total order 6000 * 15 = 90,000

Total cost = 6 + 2 + 2 = 10 * 6000 = 60,000

so income = revenue - cost

= 90,000 - 60,000 = 30,000

Note: No fixed costs are avoidable